A particle moves, so that the distance (s) in meters travel after time (t) seconds is given by s=f(t). Find the expression for velocity and acceleration, if s=cos2πt + sin2πt at t=1 second and π=180°


1
Expert's answer
2021-02-22T11:06:34-0500

Displacement of particle as a function of time is given by,

"s(t)=cos2\pi t+sin2\pi t"

Differentiating above equation with respect to time,

"\dfrac{ds}{dt}=-2\pi sin2\pi t+2\pi cos2\pi t"


"\Rightarrow v(t)=2\pi(cos2\pi t-sin2\pi t)"

We get velocity as a function of time.

Differentiating the velocity equation with respect to time,

"\dfrac{dv}{dt}=2\pi(-2\pi sin2\pi t-2\pi sin2\pi t)"

"\Rightarrow a(t)=-4\pi^{2}(sin2\pi t+cos2\pi t)"

We get acceleration as a function of time.

At t = 1 second,

"v(t=1)=2\pi(cos2\pi-sin2\pi)"

"=2\pi(1-0)"

"=2\times3.14"

"v(t=1)=6.28 ms^{-1}"


"a(t=1)=-4\pi^{2}(sin2\pi+cos2\pi)"

"= -4\pi^{2}"

"a(t=1)=-39.47ms^{-2}"